OpenStudy (anonymous):

What adds to -10 and multiplies to 6?

sammixboo (sammixboo):

?? Confusing. ANymore info?

OpenStudy (anonymous):

No this is a step to factoring :(

OpenStudy (ilovecake):

What answer did you get?

OpenStudy (ilovecake):

or did you try it yet

OpenStudy (anonymous):

The equation is 3x2 - 10x + 2 = 0, so I add the last term and first coefficient and then find factors of that number that adds to the middle term

OpenStudy (anonymous):

..... What?!?!

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

...

OpenStudy (anonymous):

So basically factors of 6 that add to -10

OpenStudy (anonymous):

(or subtract I guess)

OpenStudy (anonymous):

I outsourced because this is a tad ridiculous and I don't want to write out every possible combination of numbers that add (or add negatives) to get -10 because it's literally infinite

jimthompson5910 (jim_thompson5910):

3x^2 -10x + 2 = 0 is in the form ax^2 + bx + c = 0 where a = 3 b = -10 c = 2 plug these values into the discriminant formula d = b^2 - 4ac d = (-10)^2 - 4(3)(2) d = 100 - 24 d = 76 the discriminant is positive but it's not a perfect square. So we can't factor 3x^2 -10x + 2 (over the rationals)

OpenStudy (anonymous):

oh, I should have paid more attention then!

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

Could it also be called the quadratic formula because I can't use that again

jimthompson5910 (jim_thompson5910):

Rule: if d = b^2 - 4ac is a positive perfect square, then you can factor ax^2 + bx + c over the rationals if d = 0, then ax^2 + bx + c is a perfect square trinomial

jimthompson5910 (jim_thompson5910):

why can't you use it?

OpenStudy (anonymous):

I used it already; I can't use the quadratic formula (if that's what that is?)

OpenStudy (anonymous):

again

jimthompson5910 (jim_thompson5910):

you can use the quadratic formula to solve any quadratic equation 3x^2 -10x + 2 = 0 is a quadratic since the leading term has an x^2 in it so you plug a = 3, b = -10 and c = 2 into the quadratic formula to solve 3x^2 -10x + 2 = 0

OpenStudy (anonymous):

I know, but the question asks specifically to use a method of solving different than the one before (I used quadratic formula already)

jimthompson5910 (jim_thompson5910):

ok, you can complete the square then

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

do you know how to do so?

OpenStudy (anonymous):

I think I can do that

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

Thank you so much! Also, do you get any reward for helping people? or is this just your hobby?

jimthompson5910 (jim_thompson5910):

just a fun thing to do

OpenStudy (anonymous):

hmm cool, thanks!

jimthompson5910 (jim_thompson5910):

np